# The Gray Zone of Faithfulness: Taming Ambiguity in Unfaithfulness Detection

### v1

@@ Abstract
Qiang Ding1,2,3, Lvzhou Luo1,2,3, Yixuan Cao1,2,3*, Ping Luo1,2,3* 1Key Lab of Intelligent Information Processing of Chinese Academy of Sciences (CAS), Institute of Computing Technology, CAS, Beijing 100190, China 2State Key Lab of Al Safety, Beijing 100094, China 3University of Chinese Academy of Sciences, Beijing 100049, China {dingqiang22z,luolvzhou23s,caoyixuan,luop}@ict.ac.cn
@@ 3 Annotation Framework
The “black box” of the Su-24 jet was officially opened in Moscow on Friday in front of journalists and diplomats. … The “black box” of the Su-24 jet, shot down on November 24th, was opened in Moscow on Friday. … Explicitly-Supported
@@
A sentence $s$ is Explicitly-Supported by document $D$ if and only if i) $D\vDash s$; and ii) For each event, relation, or state in $s$, there exists a semantically equivalent event, relation, or state in $D$.
@@ 4.2 Human Annotation
Quality Assurance. The quality assurance process was conducted in multiple stages to ensure high quality. Firstly, before annotation, all annotators completed a training phase that included reviewing canonical examples and passing a quiz on challenging cases. Secondly, during annotation, annotators selected not only the final label but also intermediate decision options such as Contains Fast, Is Ambiguous, and Is Supported by Doc. The system automatically verified the consistency between these intermediate choices and the final label, and only consistent annotations could be submitted. Thirdly, after annotation, each instance underwent successive reviews by a second annotator. Fourthly, instances not classified as Ambiguous or Out-Dependent were processed by a modified version [^3] The original error detector runs four LLM-as-a-judge models, with its final prediction confirmed by a second round of LLM-as-a-judge. Our modification introduced a new LLM-as-a-judge model of GPT-5 and omitted the second round as it empirically reduced error detection recall compared to the first round alone. of the automatic annotation error detector proposed by Seo et al. (2025). Specifically, the modified detector runs five LLM-as-a-judge models (using o3-mini, GPT-4o, Gemini 2.0-Flash, Llama3.1 405B, and GPT-5), each providing their individual faithfulness predictions (Attributable, Not Attributable, or Contradictory). Potential annotation errors were flagged whenever any LLM’s prediction disagreed with the annotated label. Finally, error detection results were reviewed by the annotators to produce the final annotated dataset.
@@ 5.1 Baselines
Fine-tuned LLMs. The fine-tuned LLMs comprise Minicheck 7B [^4] https://huggingface.co/bespokelabs/Bespoke-MiniCheck-7B (Tang et al., 2024), Lynx 8B Instruct v1.1555https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-8B-Instruct-v1.1, Lynx 70B Instruct [^6] https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-70B-Instruct (Ravi et al., 2024), and ANAH-v2777https://huggingface.co/opencompass/anah-v2 (Gu et al., 2024), where Minicheck and Lynx models are binary classifiers, and ANAH-v2 model is a three-way classifier.
@@ 5.3 Results
The evaluation results for metrics except class-wise F1 are shown in Table 5.3. As shown, the most effective methods are zero-shot LLMs, specifically GPT-5 and DeepSeek-R1, which achieve balanced accuracy of $75\%\sim 80\%$, hallucination F1 of $\sim 60\%$, and ranking loss of $\sim 30\%$. Figure 4 shows the class-wise F1 of zero-shot LLMs. Across all such models, F1 values for Ambiguous and Out-Dependent remain consistently low, suggesting that LLMs have difficulty distinguishing Not Sure instances from others. We hypothesize that for Out-Dependent, this may be due to the models’ lack of access to external knowledge sources, while their parametric knowledge is insufficient to verify such cases; and for Ambiguous, this may stem from the models’ limited sensitivity to linguistic ambiguity.
Selective Prediction Evaluation. We further evaluate the zero-shot LLMs in terms of selective prediction, treating the detectors as confidence estimators (Geifman and El-Yaniv, 2017). In this setting, selective prediction assesses both the selective risk – the hallucination rate for samples whose confidence is greater than or equal to a certain threshold – and the coverage, defined as the proportion of samples with confidence greater than or equal to that threshold. Here, confidence is defined as the faithfulness degree, whose ordering has been established earlier. We vary the confidence threshold across Out-Dependent, Implicitly-Supported, and Explicitly-Supported to comprehensively evaluate the selective prediction performance. For a fixed coverage, a lower selective risk indicates better selective prediction performance.

### v2

@@ Abstract
Qiang Ding, Lvzhou Luo, Yixuan Cao*, Ping Luo* Key Lab of Intelligent Information Processing, Institute of Computing Technology, Chinese Academy of Sciences (CAS), Beijing 100190, China State Key Lab of Al Safety, Beijing 100094, China University of Chinese Academy of Sciences, Beijing 100049, China {dingqiang22z,luolvzhou23s,caoyixuan,luop}@ict.ac.cn
@@ 3 Annotation Framework
The “black box” of the Su-24 jet was officially opened in Moscow on Friday in front of journalists and diplomats. …The Su-24 was shot down by F-16 fighters on 24 November. … The “black box” of the Su-24 jet, shot down on November 24th, was opened in Moscow on Friday. … Explicitly-Supported
@@
A sentence $s$ is Explicitly-Supported by document $D$ if and only if i) $s$ is logically implied by $D$, or $D\vDash s$; and ii) For each event, relation, or state in $s$, there exists a semantically equivalent event, relation, or state in $D$.
@@ 4.2 Human Annotation
Quality Assurance. The quality assurance process was conducted in multiple stages to ensure high quality. Firstly, before annotation, all annotators completed a training phase that included reviewing canonical examples and passing a quiz on challenging cases. Secondly, during annotation, annotators selected not only the label but also intermediate decision options such as Contains Fast, Is Ambiguous, and Is Supported by Doc. The system automatically verified the consistency between these intermediate choices and the final label, and only consistent annotations could be submitted. Thirdly, after annotation, each instance underwent successive reviews by a second annotator. Fourthly, instances not classified as Ambiguous or Out-Dependent were processed by a modified version [^3] The original error detector runs four LLM-as-a-judge models, with its final prediction confirmed by a second round of LLM-as-a-judge. Our modification introduced a new LLM-as-a-judge model of GPT-5 and omitted the second round as it empirically reduced error detection recall compared to the first round alone. of the automatic annotation error detector proposed by Seo et al. (2025). Specifically, the modified detector runs five LLM-as-a-judge models (using o3-mini, GPT-4o, Gemini 2.0-Flash, Llama3.1 405B, and GPT-5), each providing their individual faithfulness predictions (Attributable, Not Attributable, or Contradictory). Potential annotation errors were flagged whenever any LLM’s prediction disagreed with the annotated label. Finally, error detection results were reviewed by the annotators to produce the final annotations.
@@ 5.1 Baselines
Fine-tuned LLMs. The fine-tuned LLMs comprise Minicheck 7B [^4] https://huggingface.co/bespokelabs/Bespoke-MiniCheck-7B (Tang et al., 2024), Lynx 8B Instruct v1.1555https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-8B-Instruct-v1.1, Lynx 70B Instruct [^6] https://huggingface.co/PatronusAI/Llama-3-Patronus-Lynx-70B-Instruct (Ravi et al., 2024), and ANAH-v2777https://huggingface.co/opencompass/anah-v2 (Gu et al., 2024), where Minicheck and Lynx models are binary classifiers, and ANAH-v2 is a three-way classifier.
@@ 5.3 Results
The evaluation results for metrics except class-wise F1 are shown in Table 5.3. As shown, the most effective methods are zero-shot LLMs, specifically GPT-5 and DeepSeek-R1, which achieve balanced accuracy of $75\%\sim 80\%$, hallucination F1 of $\sim 60\%$, and ranking loss of $\sim 30\%$. Figure 4 shows the class-wise F1 of zero-shot LLMs. The results reveal that DeepSeek-R1’s better hallucination F1 compared to GPT-5 (shown in Table 5.3) can be attributed to its higher F1 score in the Fabricated class, as this class accounts for the majority of the Unfaithful classes. Across all such models, F1 values for Ambiguous and Out-Dependent remain consistently low, suggesting that LLMs have difficulty distinguishing Not Sure instances from others. We hypothesize that for Out-Dependent, this may be due to the models’ lack of access to external knowledge sources, while their parametric knowledge is insufficient to verify such cases; and for Ambiguous, this may stem from the models’ limited sensitivity to linguistic ambiguity.
Selective Prediction Evaluation. To investigate whether the Explicitly-Supported prediction serves as a reliable indicator of faithfulness confidence, we frame zero-shot LLM detectors as confidence estimators and assess their performance under the framework of selective prediction (Geifman and El-Yaniv, 2017). In this setting, selective prediction assesses both the selective risk – the hallucination rate for samples whose confidence is greater than or equal to a certain threshold – and the coverage, defined as the proportion of samples with confidence greater than or equal to that threshold. Here, confidence is defined as the faithfulness degree, whose ordering has been established earlier. We vary the confidence threshold across Out-Dependent, Implicitly-Supported, and Explicitly-Supported to comprehensively evaluate the selective prediction performance. For a fixed coverage, a lower selective risk indicates better selective prediction performance.